Three boys are going to a party and comparing how much money they have. sam works out that he has $6 more than Mark and $12 more than Dean. If they have $75 altogether, how much does Sam have?
Answer by dolly(163)

Let Sam have $x initially
==>
Mark has $(x-6)
and
Dean has $(x-12)
As the total amount is $75
x+(x-6)+(x-12)=75
==> 3x-6-12=75
==> 3x-18=75
Adding 18 on both sides,
==> 3x-18+18=75+18
==> 3x=93
==> 3x/3=93/3
==> x=31
==> Sam has $31